MONT 105N – Analyzing Environmental Data Paper 2 Assignment – What Should A College Education Teach You? due: April 15, 2019

Two options

For this writing assignment you have two options to choose from:

1. The so-called “Green New Deal” proposed by the recently-elected Congressional representative Alexandria Ocasio-Cortez from New York together with other Democrats has been in the news since it was submitted as a draft bill in Congress in February. It has attracted significant criticism from Republicans and more conservative Democrats for several different sets of reasons. But what exactly does it say? This topic will require some research. You will find the actual text of the draft legislation at:

 https://www.congress.gov/116/bills/hres109/BILLS-116hres109ih.pdf

In addition to reading this and figuring out what is proposed, please find at least two opinion pieces from reputable publications that take opposite sides about this proposal and summarize their arguments. (You might start with major newspapers like the New York Times, Washington Post, Boston Globe, and magazines like The Atlantic, Harpers, etc. Articles or op-ed pieces from their websites are OK.) You can add your own opinions on the issues raised here in a concluding paragraph, but the main point of this will be to learn exactly what the “Green New Deal” is proposing and why certain people have supported it, while others have been so scornful about it. For this option, you should provide a listing of sources you consulted and identify any direct quotations or paraphrases taken from the sources you consulted using footor endnotes.

2. The second option is based on a draft list of learning outcomes that has been circulated by a faculty committee that will be developing a replacement for the current system of Common Area Requirements. This is available from a link on the course homepage (right next to the link where you downloaded this assignment). It seems that this committee has been working with only limited student input so far (although there are two student representatives on the committee that drafted this statement). So I’m interested in what you would think about what they have come up with at this point. Note: When you read the statement, you will see that it is nowhere near a specific policy for course requirements. Instead, this (or the eventual version of this that will come out of a lot of discussions that are starting about now) is like the blueprint for developing a new system of specific course requirements. You should compose an essay of about 3 pages (double-spaced, one side) addressing some or all of the following questions:

1. What are your first reactions to this statement? Recall, this is what some faculty think you should be getting from your education here. How does it relate to how you think about your education? Are these sorts of learning outcomes anything you have ever considered?

2. Is this too lenient or “easy?” Is it reasonable? Is it too much to expect?

3. Are these the right learning outcomes to be guiding everyone’s educational experience? Or are they undesirably slanted in some way? If you think there are issues here, what else should the faculty working on this issue be thinking about? What’s missing or not present that should be there? Are there things here that we should not be focusing on?

4. This one might be hard at your current stage in your educations, but here goes anyway: What should a system of common requirements that all students have to satisfy look like? Should that take into account the special history and mission of Holy Cross as a Catholic, Jesuit college? (For example, should everyone have to take a year of Latin? That’s probably unrealistic, but it’s a great way to solidify one’s grasp of issues of grammar, English vocabulary, etc.) What are the things all students should know or be able to do (content knowledge, skills, “habits of mind” or whatever)?

Note: Unlike the first option, most of these questions are asking for personal opinions. So it is not expected or necessary that you will consult any outside sources if you choose this option. Still, you should try as much as possible to justify what you are saying by citing details from your own experiences so far.

Important Note: It is perfectly OK with me if you want to share a rough draft of your essay with other students in the class, or with me, to get comments and suggestions. “Testing out” your ideas that way is often an important part of getting them down on paper in an effective form.

Prepare your essay using Google Docs, Microsoft Word or similar software and submit as a file attachment to this email address:

                   jlittle@holycross.edu

Your assignment should arrive in my inbox no later than 5:00pm on Monday, April 15. (Note: This date makes me think of income tax day for some reason! If any of you are preparing Federal and Massachusetts state income tax returns, you have until April 17 to submit because of the Patriot’s Day holiday in Massachusetts.) Aim for around 3 doublespaced pages. (Going a bit short or long is OK, though, as long as you have addressed one of the options above adequately and succinctly.)